# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v1

@@ 6 Experiments
We conduct a preliminary empirical comparison between OPO and GSPO on mathematical reasoning tasks using the VERL framework.
@@ 6.1 Experimental Setup
We use Qwen3-1.7B as the base model and train on MATH Level 3 problems. Both methods use identical training configurations: 4 epochs, batch size 32, learning rate $2\times 10^{-6}$, and 6 rollout generations per prompt.
Configurations.
  • OPO: $\alpha=0.6$, $\mu=1.0$, adaptive $\tau$ with range $[0.2,1.5]$.
  • GSPO: Standard sentence-level GRPO with advantage normalization.
@@ 6.2 Results
Figure 1 shows the training accuracy (mean reward) over 224 training steps. Both methods exhibit similar convergence behavior, with OPO achieving slightly higher final accuracy (approximately 76% vs 71% averaged over the last 20 steps). The difference is modest but consistent across training.
Figure 1: Training dynamics comparing OPO ($\alpha=0.6$, $\mu=1.0$) and GSPO on Qwen3-1.7B Math RL. Solid lines show 7-step moving averages; faded lines show raw values. Both methods achieve comparable final performance, with OPO showing a slight edge.
@@
Figure 2 illustrates the gradient norm dynamics across training. OPO exhibits consistently higher gradient norms, suggesting more active optimization signal in accordance with the theoretical prediction of non-saturating linear gradients. The GSPO baseline shows lower and more stable gradient norms, potentially reflecting the implicit regularization from advantage normalization.
Figure 2: Gradient norm comparison. OPO maintains higher gradient norms throughout training, consistent with the theoretical prediction of non-saturating gradient dynamics.
These preliminary results suggest that OPO achieves comparable or slightly better performance than GSPO on this task, while exhibiting gradient dynamics that align with theoretical predictions. Both methods converge stably without training instabilities. We emphasize that these experiments are preliminary and conducted on a single model scale.
@@ 7 Discussion
The OPO objective (Equation˜6) involves expectations under $\pi_{\mathrm{ref}}$. In practice:

### v2

@@ 6 Experiments
We conduct an empirical comparison of OPO against three strong baselines—GRPO, GSPO, and DAPO—on mathematical reasoning tasks using the VERL framework.
@@ 6.1 Experimental Setup
We use Qwen3-1.7B as the base model and train on MATH Level 3 problems. All methods use identical training configurations: 4 epochs, batch size 32, learning rate $2\times 10^{-6}$, and 6 rollout generations per prompt.
Baseline Configurations.
  • GRPO (Group Relative Policy Optimization) [4]: Standard token-level policy gradient with group-normalized advantages. Uses vanilla PPO loss mode without a value function critic.
  • GSPO (Group Sentence-level Policy Optimization): A variant of GRPO that applies advantage normalization at the sentence/step level rather than the token level, aiming for more granular credit assignment.
  • DAPO (Dynamic Advantage Policy Optimization): An advanced baseline incorporating four key mechanisms:
    1. 1.
      Clip Higher: Raises the PPO clip upper bound (e.g., $1+\epsilon\to 1.28$) to mitigate entropy collapse and preserve exploration.
    2. 2.
      Dynamic Sampling: dynamically filters "all-correct" or "all-wrong" samples that contribute zero advantage, focusing training on informative samples.
    3. 3.
      Token-Level Policy Gradient: Standardizes advantage weights across tokens in a mini-batch to prevent gradient dilution in long sequences.
    4. 4.
      Overlong Reward Shaping: Applies a length-aware penalty and masks gradients for truncated responses to reduce noise.
  • OPO: The proposed method with $\alpha=0.6$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$.
@@ 6.2 Results
Overall Performance.
Table 1 summarizes the key metrics across all four algorithms.
Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3. Reward and gradient norm are averaged over the final 20 training steps.
Algorithm Mean Reward Grad Norm Characteristics
GRPO 0.686 0.61 Standard baseline, stable
GSPO 0.713 0.50 Strong, some gradient decay
DAPO 0.67* 0.22 Conservative, early plateau
OPO (Ours) 0.756 0.90 Best reward, non-saturating gradients
Algorithm Characteristics.
  • GRPO: Serves as the canonical baseline. Achieves reasonable performance (69% acc) with moderate gradient norms. The token-level normalization provides stability but may average out strong signals.
  • GSPO: Improves upon GRPO (71% acc) via sentence-level credit assignment. However, gradient norms (0.50) are lower than OPO’s, suggesting some saturation in high-confidence regimes.
  • DAPO: Exhibits the lowest gradient norms (0.22) and plateaus early. The conservative clipping constraints appear overly restrictive for reasoning tasks where continued learning is beneficial. *Note: Reward fluctuated between 0.59–0.67.
  • OPO: Achieves the highest mean reward (0.756) while maintaining the largest gradient norms (0.90). This empirically validates the theoretical prediction: the $\chi^{2}$ geometry provides non-saturating linear gradients, allowing the model to continue learning where others plateau.
@@
Figure 3 shows the training accuracy (mean reward) over 224 training steps.
  • Early Stage: GRPO exhibits the fastest initial learning, surpassing OPO in the first 50 steps. This is consistent with standard policy gradient methods having aggressive early updates.
  • Late Stage Cross-over: As training progresses (after step 100), GRPO and GSPO begin to saturate. In contrast, OPO maintains a steady improvement rate, eventually overtaking all baselines to reach the highest final accuracy ($\sim$76%). This confirms that the $\chi^{2}$-induced linear drive prevents the “vanishing gradient” problem common in KL-constrained methods as confidence increases.
Figure 3: Training dynamics on Qwen3-1.7B Math RL. Note the cross-over behavior: GRPO (green) starts strong but plateaus, while OPO (red) sustains improvement and achieves the highest final performance.
@@
Figure 4 illustrates the gradient norm dynamics across training. OPO exhibits consistently higher gradient norms, validating the non-saturating property of the $\chi^{2}$ geometry. DAPO shows severely diminished gradients, explaining its early plateau.
Figure 4: Gradient norm comparison. OPO maintains healthy gradient norms (0.9) throughout training, while DAPO (0.22) and GSPO (0.50) exhibit gradient decay.
Entropy Dynamics.
Figure 5 depicts the policy entropy throughout training. OPO maintains consistently higher entropy levels compared to baselines. We attribute this to the granularity of supervision: GRPO and DAPO utilize token-level updates, imposing dense supervision that forces the policy to collapse rapidly into specific phrasings. In contrast, OPO (and GSPO) employs sequence-level updates. This “sparse” supervision aligns the total trajectory probability without micromanaging individual tokens, thereby preserving policy diversity and exploration potential.
Figure 5: Entropy dynamics. OPO maintains higher entropy than baselines, preventing premature mode collapse and enabling sustained exploration throughout the training process.
Summary of Findings.
The experimental results confirm that OPO’s orthogonalized design—combining $\alpha$-weighted sampling with $\chi^{2}$ optimization geometry—prevents gradient saturation and enables continued learning in high-confidence regimes.
@@ 7 Discussion
The OPO objective (Equation˜7) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
@@ 8 Conclusion
Appendix A Theoretical Proofs and Derivations
A.1 Log-Ratio Approximation Error
Lemma A.1 (Log-Ratio Approximation Error).
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation. Let $\delta=\|\Delta_{\theta}\|_{\infty}$. For $\delta<1$ (trust region regime):
$|v_{\theta}-\Delta_{\theta}|\leq\frac{1}{2}\delta^{2}e^{\delta},\quad|v_{\theta}^{2}-\Delta_{\theta}^{2}|\leq\delta^{3}e^{2\delta}$ (18)
Consequently, the difference between the exact ratio loss and the log-approximate loss scales as $O(\mathbb{E}[\Delta_{\theta}^{3}])$.
Proof.
Using Taylor expansion $e^{x}=1+x+\frac{x^{2}}{2}+\dots$, we have $v=\Delta+\frac{\Delta^{2}}{2}+O(\Delta^{3})$. The bounds follow from standard remainder estimation for the exponential function. ∎
A.2 Equivalence to $\chi^{2}$-Constrained Maximization
Proposition A.2 (Lagrange Dual Equivalence).
Consider the problem of maximizing alignment with the target $\omega_{\alpha}$ subject to a functional $\chi^{2}$ trust region:
$\max_{v\in L^{2}(\pi_{\mathrm{ref}})}\mathbb{E}_{\pi_{\mathrm{ref}}}[\omega_{\alpha}(y)v(y)]\quad\text{s.t.}\quad\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]\leq\epsilon$ (19)
The Lagrangian relaxation of this problem is exactly the OPO objective $\mathcal{L}_{\text{OPO}}=-\mathbb{E}[\omega_{\alpha}v]+\frac{\mu}{2}\mathbb{E}[v^{2}]$, where the regularization coefficient $\mu$ is the Lagrange multiplier corresponding to the trust region radius $\epsilon$.
The optimal solution is $v^{*}(y)=\frac{1}{\mu}\omega_{\alpha}(y)$. This duality implies that $\alpha$ purely shapes the objective (alignment direction), while $\mu$ purely enforces the feasibility radius (optimization geometry).
A.3 Distributional Stability Guarantees
Proposition A.3 ($\chi^{2}$ Controls Total Variation).
Let $TV(\pi,\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[|t(y)-1|]=\frac{1}{2}\mathbb{E}[|v(y)|]$. By Jensen’s inequality:
$TV(\pi,\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}[|v|]\leq\frac{1}{2}\sqrt{\mathbb{E}[v^{2}]}$ (20)
Thus, bounding the $\chi^{2}$ norm $\mathbb{E}[v^{2}]\leq\epsilon$ guarantees that the policy remains within a $\sqrt{\epsilon}/2$-radius of the reference in Total Variation distance.